You are evaluating the balance sheet for SophieLex’s Corporation. From the balance sheet you find the following balances: cash and marketable securities = $480,000; accounts receivable = $1,040,000; inventory = $1,940,000; accrued wages and taxes = $420,000; accounts payable = $720,000; and notes payable = $440,000. Calculate SophieLex’s current ratio. (Round your answer to 2 decimal places.) Calculate SophieLex’s quick ratio. (Round your answer to 2 decimal places.)
Answer #1

Ans Current Ratio = 2.19

Quick Ratio = 0.96

Current Assets = Cash & Marketable Securities + Accounts receivable + Inventory

                         = 480000 + 1040000 + 1940000

                          = 3460000

Current Liabilities = Accrued wages and taxes + Accounts payable + notes payable

                       = 420000 + 720000 + 440000

                       = 1580000

Current Ratio = Current Assets/ Current Liabilities

                        = 3460000 / 1580000

                       = 2.19 : 1

Quick Ratio = ( Current Assets - Inventory) / Current Liabilities

                    = (3460000 - 1940000) / 1580000

                   = 0.96 : 1